OpenStudy (wild_child117):

Solve for x. 3/8 (2x+16) − 2 = 13 Enter your answer in the box. x =

OpenStudy (wild_child117):

@zepdrix

OpenStudy (quickstudent):

Here's how you solve for x: 3/8 (2x + 16) − 2 = 13 0.375(2x + 16) - 2 = 13 (0.375 * 2x) + (0.375 * 16) - 2 = 13 0.75x + 6 - 2 = 13 0.75x + 6 - 2 + 2 = 13 + 2 0.75x + 6 = 15 0.75x + 6 - 6 = 15 - 6 0.75x = 9 0.75x/0.75 = 9/0.75 x = 12

OpenStudy (wild_child117):

thank you

OpenStudy (sapphiremoon):

Okay, the first thing to do is to get the x term by itself. So we'll add 2 to both sides to get: \[\frac{3}{8}(2x +16) = 13 + 2 =15\] That x still has some stuff attached to it, let's get that 3/8 off by dividing both sides by 3/8: \[(2x +16) = 15 \div \frac{3}{8} = 15 \times \frac{8}{3} = \frac{120}{3} = 40\] Now we can move the 16: \[2x = 40 -16 = 24\] And divide both sides by 2: \[x = 40 \div 2 = 12\] Make sense? I finished it cause I'd already done most of the work, lol.

OpenStudy (wild_child117):

Solve for x. x + 3.2 = −3.5 Enter your answer in the box. x =

OpenStudy (wild_child117):

@zepdrix

OpenStudy (wild_child117):

@quickstudent

OpenStudy (sapphiremoon):

Just subtract 3.2 from both sides. You'll get a negative number with a magnitude bigger than -3.5.